OpenStudy (anonymous):

A cylindrical water glass has a radius of 9 centimeters and a height of 12 centimeters. What is its volume rounded to the nearest hundredth?

OpenStudy (anonymous):

answer in cm cubed if you can :D

OpenStudy (rulnick):

pi r^2 h = pi (9 cm)^2 (12 cm) = pi (81) (12) cm^3

OpenStudy (rulnick):

... or about 3054 cm^3.

OpenStudy (anonymous):

thanks!

OpenStudy (rulnick):

welcome